OpenStudy (zenmo):

Find the exact area of the surface obtained by rotating the curve about the y-axis.

OpenStudy (zenmo):

\[x = \frac{ 1 }{ 3 }(y^2+2)^{3/2}, 1 \le y \le 2\]

OpenStudy (zenmo):

\[S = \int\limits_{a}^{b}2 \pi y ds \]